# Data Assetization via Resources-decoupled Federated Learning

### v1

@@ 3.2. Tripartite Stackelberg Model
In this section, we specify the strategies of $S$, $D$, and $C$, each trying to maximize its profit. We model the Stackelberg game for three parties and define the Stackelberg-Nash equilibrium to achieve global utility optimization.
The strategy profile denoted as $\left\langle\eta^{*},Q^{*},G^{*}\right\rangle$, consists of the strategies of the model owner $S$, data owners $D_{n}(n=1,\ldots,N)$, and computing centers $C_{m}(m=1,\ldots,M)$. These strategies are taken in order: firstly, the model owner $S$ determines the optimal total model payment $\eta^{*}$; then, each data owner $D_{n}$ decides the profitable quantity and quality of data to provide, i.e., the model quality contribution $q_{n}^{*}$; finally, each computing center $C_{m}$ determines the beneficial matching relationship with specific data owner, i.e. $G^{*}$ based on the optimal quantity of data undertaking $d_{m}$. To protect data privacy, we perform a one-time matching between data owners and the computing centers. Therefore, we formalize the resource-decoupled FL strategy as a Tripartite Stackelberg Model in Def. 1, where the model owner acts as the leader, data owners act as secondary leaders, and the computing centers act as followers. The Stackelberg model establishes the sequential achievement of sub-utility optimization goals based on the major-minor relationships between the three parties. Within this framework, $N$ data owners and $M$ computing centers adopt strategies in their inner Nash game, ultimately achieving global utility optimization.
@@
The game comprises the strategies of the model owner $S$, the data owners $D_{i}(i=1,\ldots,N)$, and the computing centers $C_{i}(i=1,\ldots,M)$ to achieve global utility optimization in resource-decoupled FL.
(8) $\left\{\begin{array}[]{ll}\text{Leader }S:\quad\eta^{*}=\arg\max\limits_{\eta}U_{s},\\[5.0pt] \text{Sub-Leaders }D_{n}:\quad q_{n}^{*}=\arg\max\limits_{q_{n}}U_{n},&n=1,\dots,N,\\[5.0pt] \text{Follows }C_{m}:\quad G^{*}=\arg\max\limits_{G}U_{m},&m=1,\dots,M.\end{array}\right.$
Our goal is to generate the strategy profile $\left\langle\eta^{*},Q^{*},G^{*}\right\rangle$ of the tripartite Stackelberg model defined by Equ. 8 to achieve the SNE. Specifically, for the sub-game between data owners, we derive its corresponding Nash equilibrium (Appendix A.1), and the decisions of the three parties collectively achieve the SNE. We define the Stackelberg-Nash equilibrium in the resource-decoupled FL, as shown in Equ. 9.
@@
(9) $\left\{\begin{array}[]{ll}U_{s}(\eta^{*},Q^{*},G^{*})\geq U_{s}(\eta,Q^{*},G^{*}),\\[8.0pt] U_{n}(\eta^{*},q_{n}^{*},q_{-n}^{*},G^{*})\geq U_{n}(\eta^{*},q_{n},q_{-n}^{*},G^{*}),&n=1\ldots N\\[8.0pt] U_{m}(\eta^{*},Q^{*},G^{*})\geq U_{m}(\eta^{*},Q^{*},G),&m=1\ldots M\end{array}\right.$
where $Q_{n}^{*}=\{q^{*}_{1},q^{*}_{2},...,q^{*}_{N}\}$ is the Nash equilibrium solutions of the sub-games. SNE indicates that each of the three parties takes its optimal strategy, which maximizes its utility. No one can add its own utility by unilaterally changing its strategy.
@@ 3.3. Analysis of Tripartite Stackelberg Model
The policy space of the computing center is the quantity of data undertaken. This interval is $[0,\max|d_{m}|]$, which is bounded, closed, and continuous. Based on computing the utility function $U_{m}$ of the center, we obtain the first and second derivatives of $U_{m}$.
@@
Subsequently, we set the first derivative in Equ.10 to zero to obtain the optimal strategy for the computing center.
(12) $d_{m}^{*}=\sqrt{\frac{\lambda\sum_{i\neq m}d_{i}\sum_{n=1}^{N}\rho x_{n}}{\varepsilon\sigma_{m}}}-\sum_{i\neq m}d_{i}$
@@
For data owner $q_{n}$, there is a unique Nash equilibrium optimal strategy $q_{n}^{*}$ to maximize its utility $U_{n}$ for the given $\eta$ and $q_{-n}$.
@@
The policy space for the data owner is to decide how much model quality to contribute. This interval is $[0,|Xn|]$, which is bounded, closed, and continuous. According to the utility function $U_{n}$ of the data owner, we obtain the first and second order derivatives of $U_{n}$ as shown in Equ. 13 and Equ. 14.
@@
For the model owner, there is a unique optimal $\eta^{*}$ to optimize the utility $U_{s}$ for given the parameter of $\alpha$ and objective data quality provided by the data owner $f_{n},n\in(0,N]$.
@@
The range of $\eta$ is $[0,\infty]$ for each model owner. Thus, this set of strategies is clearly convex and continuous. Although the value of $\eta$ can be infinite in theory, practically the model owner will not give infinite total payment of the model, and there is also the effect of diminishing marginal utility. This leads to the fact that $\eta$ is necessarily bounded and closed. We obtain the first and second order derivatives as shown in equ. 16 and Equ. 17.
@@
where $T_{n}=\frac{(N-1)}{\lambda\rho\sum_{i=1}^{N}{\frac{1}{f_{i}}}}(1-\frac{N-1}{f_{n}\sum_{i=1}^{N}{\frac{1}{f_{i}}}})$, Since $g(\cdot)$ is a convex function, and to simulate the diminishing marginal benefit, here we adopt the ln function: $g(x)=ln(1+x)$. ∎
@@
The existence and uniqueness of SNE can be deduced by the property that the strategy space is a convex and compact subspace of the Euclidean space while the utility functions are concave. In the case of the model owner, it can be justified that the optimal utility $U_{s}$ is obtained only at $\eta^{*}$ derived by direct derivation due to the strictly concave property of Equ. 7, leading to the first inequality in Def. 2 holding uniquely at $\eta^{*}$. Similarly, a unique Nash equilibrium can be justified, and the second inequality holds for every data owner at $q_{n}^{*}$. However, for privacy reasons, we don’t split the data in the computing center; instead, we use the Gale-Shapley algorithm to achieve one-time matching with the data owner according to the optimal $d_{m}^{*}$. According to (Teo et al., 2001; Gokhale et al., 2024), the Gale-Shapley algorithm satisfies the relaxed Nash equilibrium, thus satisfying the third inequality. Therefore, SNE exists since the set of inequalities in Equ. 9 can be satisfied at $\left\langle\eta^{*},Q^{*},G^{*}\right\rangle$. ∎
@@ 3.4. QD-RDFL
In this section, we design a Quality-aware Dynamic Resource-decoupled FL algorithm to identify the optimal strategy profile, in which a dynamic optimization mechanism is designed to improve the optimal strategy by evaluating the model contribution of data owners.
Using the backward induction method, we first calculate the optimal strategy of the model owner $\eta$ according to Equ. 18. We then calculate the optimal strategy $Q_{n}$ for data owner $D_{n}$ according to Equ.14. Since $Q_{n}$ is evaluated along two dimensions, data quality is a posterior metric assessed by the server after training, and data quantity $D_{n}$ is a prior decision determined immediately. Therefore, we initialize $D_{n}$ with an initial $f_{n}$ for computation and dynamically adjust $f_{n}$ based on the real training results. Finally, we calculate the optimal $d_{m}$ according to Equ. 12, and construct the optimal matching $G^{*}$ using the Gale-Shapley algorithm (Teo et al., 2001; Gokhale et al., 2024). The matching remains unchanged during subsequent computations for private protection, meaning that changes in $f_{n}$ will not affect the matching relationship.
We evaluate the quality of data based on its genuine contribution to ML. The most accurate method is to evaluate each local model by computing its test accuracy. However, this approach incurs additional costs and resource consumption. In contrast, accessing training loss does not require extra overhead, as it is generated during training. Therefore, we use the reduction in training loss during each iteration to evaluate $f_{n}$ as (Hu et al., 2024) and shown in Equ. 19.
@@
where $\text{loss}_{t}$ denotes the loss in a specific round $t$ , $t_{s}$ represents the start time of validation, and $t_{e}$ represents the end time. Clients must submit their local model updates within the specified time range; otherwise, the edge server will reject the submission.
@@
0:  $D=\{D_{1},D_{2},...,D_{N}\}$ , $C=\{C_{1},C_{2},...,C_{M}\}$, $\rho$, $\varepsilon$, $\alpha$, buffer $F=\{F_{1},F_{2},...,F_{M}\}$
0:  $\eta^{*}$ for model owner, $x^{*}$ for each data owner, matching $G$ for data owner and computing center
1:  Calculate $\eta^{*}$
2:  Calculate $x^{*}$ for each data owner
3:  $G=Gale-Shapley(D,C)$
4:  for each $D_{n},n\in[1,N]$ do
5:     Randomly pick $x_{n}^{*}$ date pieces from its dataset
6:     Send the data according to $G$
7:  end for
8:  for each round $t=1,2,...$ do
9:     for each $C_{m},m\in[1,M]$ do
10:        $w^{t+1}_{m},f^{t+1}_{m}\leftarrow\text{ClientUpdate}(m,w^{t})$
11:        Add $f^{t+1}_{m}$ to buffer $F_{m}$ and calculate $f_{m}$
@@
13:     Normalized each $f_{m}$
14:     Recalculate $\eta^{*}_{t+1}$ and $x^{*}_{t+1}$ for each data owner
@@
16:        The data owner pays extra
@@
18:     $w^{t+1}\leftarrow\frac{1}{M}\sum^{M}_{m=1}w^{t+1}_{m}$
19:  end for
@@
0:  computing center $m$, global model parameter $w$
0:  local model parameter $w_{m}$, data quality $f$
@@
6:  $f=loss(t_{s})-loss(t_{e})$
7:  return  $w_{m}$, $f$
Consequently, the methodology steps for data assetization via resource-decoupled FL are outlined below and illustrated in Fig. 1. The QD-RDFL algorithm is presented in Alg. 1 and Alg. 2.
@@
  • Step 2: Initial Optimal Strategy. $S$, $D$, and $C$ compute the optimal strategy profile $\langle\eta^{*},Q^{*},d^{*}\rangle$ for the first time based on Equ. 8 and 9.
  • Step 3: Gale-Shapley Matching Data and Computing Resources. $D$ is matched with $C$ based on the initial optimal strategy, ensuring secure data transmission.
@@
We evaluate the quality of data based on its real contribution to ML. The most accurate method is to evaluate each local model by computing its test accuracy. However, this approach incurs additional costs and resource consumption. In contrast, accessing training loss does not require extra overhead, as it is generated during training. Therefore, we use the reduction in training loss during each iteration to evaluate $f_{n}$ as (Hu et al., 2024) and shown in Equ. 19.
In Step 5, after the data quality $f_{n}$ is adjusted according to the training results, the model owner and data owners sequentially adjust the total model payment and the amount of data provided. Subsequently, data owners providing more data pay additional training fees to their corresponding computing centers.
@@ 4. Experiments
Q4: How do different components (i.e., dynamic adjustment mechanisms) affect the global model performance?
@@ 4.1. Experiments Setting
We conduct extensive experiments on various vision datasets, including MNIST (LeCun et al., 1989), CIFAR-10 (Krizhevsky, 2009), and CIFAR-100 (Krizhevsky, 2009). Our training utilizes a widely used and well-established model architecture: a Convolutional Neural Network (CNN). For the simpler MNIST dataset, we use CNN, which consists of two convolutional layers and two fully connected layers. For datasets CIFAR-10 and CIFAR-100, we use CNN composed of three convolutional layers and four fully connected layers. The FL training process involves 100 iterations, including three local epochs of client training. Cross-entropy loss is employed for all classification tasks. Meanwhile, we adopt the Adam optimizer with a learning rate of 0.001. All experiments are conducted on PyTorch 1.11.0 and RTX 4090.
To simulate the variation in data quality provided by different data owners, we add Gaussian noise with varying intensity to the data from different owners. In addition, considering that data quality is difficult to measure accurately, we calculate the mean square error (MSE) of the images after and before noise addition and use 1-MSE to represent the initial data quality of the data owner so as to ensure that it is as reasonable as possible. Furthermore, in order to represent the relative position of $f_{n}$, we normalize its value. We evaluate the data quality based on its real contribution to ML using the reduction in training loss during each iteration to evaluate $f_{n}$ according to Equ. 19. We then performance a normalization of the values to measure the relative position.
@@ 4.2. Experimental Results
SNE verification (RQ1). To confirm the existence and uniqueness of the SNE of QD-RDFL, we set $N=10$ and observe the variation of utility under different strategies of the model owner and the data owner on the MNIST dataset, respectively.
@@
Figure 3. Effect of initial reported $f_{n}$ to $U_{s}$ and the accuracy of gambol model, (a) comparison of $U_{s}$ with different proportions of data owners and different report deviation ratio $f_{n}$, and (b) the accuracy of gambol model with different $f_{n}$.
As shown in Fig. 2a, the quantity and strategies remain constant, and the curve of the model owner’s utility with the increasing total model payment is convex, with its highest point representing the optimal response. The optimal payment is 1.33 at the Stackelberg equilibrium, and the corresponding utility is about 0.57. The uniqueness of the Nash equilibrium among the data owners is demonstrated in Fig. 2a. For the selected data owner, we retain the strategy of the other data owners and the payment of the model owner and then vary its contribution to the quality of the model. The results show that the utility curve of the data owner is convex. And the maximum utility of 0.0066 is obtained when the model quality contribution is 0.087. From this, we conclude that each data owner can maximize their utility by determining a strategy for allocating computing resources.
@@
Figure 6. The gambol model performance achieved by QD-RDFL in each round on different datasets
Effect of initial $f_{n}$ (RQ2). To observe the effect of the initial $f_{n}$ on the utility of the data owner and the utility of the model owner, we design the following two sets of experiments. For data owner: we select several data owners for the experiment. For each chosen data owner, we fix the $f_{n}$ values of the other data owners and increase or decrease the selected data owner’s $f_{n}$. We then observe the variation in their utility. For model owner: we increase or decrease the given data owner’s $f_{n}$ and observe the resulting variation in the model owner’s utility.
For the specific selected data owners, each of one’s utility $U_{n}$ increases with the report deviation of $f_{n}$ on MNIST and CIFAR-10 datasets, as shown in Appendix A.2, Fig.LABEL:fig-7. The results suggest that the utility of individuals in data centers is sensitive to data quality contributions. Therefore, using data quality contributions based on real training as a criterion for distributing profits is reasonable and necessary since it prevents other data owners from getting the utility they deserve.
@@
Effect of initial $f_{n}$ (RQ3). To verify the incentive effect of our QD-RDFL, we perform comparative experiments with the following baselines on different datasets.
@@
  • Random $\eta$: The model owner publishes an FL training task with a random payment value, and all users use this payment to determine their strategies.
@@
Varying numbers of data owners, we observe the utility of model owners $U_{s}$ and the utility of data owners $U_{n}$, respectively. As shown in Figs. 4a, 4b and 4c, on different datasets as $N$ increases, more data owners participate in FL training, providing more contributions to model quality. However, due to the diminishing marginal effect, the increased speed of the model benefit is slow. The utility of the model owner $U_{s}$ with different $N$ our QD-RDFL is always optimal compared to the benchmarks. The experimental results demonstrate that QD-RDFL can achieve optimal utility and effective courage data owners in different configurations.
Figs. 5a, 5b and 5c show the average of $U_{n}$ with different $N$ on the different datasets. As $N$ increases, the average utility of the data owner decreases. This is because as $N$ increases, more data owners participate in FL training while the overall payment of the model grows relatively slowly. Therefore, the average utility of the data owner becomes smaller. As shown, the average utility of our QD-RDFL always outperforms compared with Random and Fixed mechanisms. The experimental results prove that data owners get better utility, so we also effectively incentivize data owners. It is worth noting that since each group of the experiment runs, a new value of $\eta$ is randomly generated so that Random always oscillates.
@@
Ablation study (RQ4). To evaluate the effectiveness of the proposed dynamic adjustment mechanism, we set $N=10$ and design the following ablation experiment, in which QD-RDFL without Dynamic Adjustment is set. The model owner and the data owners first determine initial strategies based solely on the data quality that the data owners provide. During the later stages of training, we dynamically adjust the data quality in different phases and then determine new strategies.
In Fig.6, by comparing the model accuracy variance between the two groups over 70 training iterations, we find that on the three datasets, the model accuracy of QD-RDFL with dynamic adjustment mechanism is better than that of removing this module in 70 iterations. This is because the dynamic adjustment module can provide a more accurate measure of data quality, so that data owners with better data quality can contribute more of their own data and make the trained model better.
5. Conclusion and Future works

### v2

@@ 3.2. Tripartite Stackelberg Model
In this section, we specify the strategies of $S$, $D$, and $C$, each trying to maximize its profit. We model the Stackelberg game for three parties and define the SNE to achieve global utility optimization.
The strategy profile denoted as $\left\langle\eta^{*},Q^{*},G^{*}\right\rangle$, consists of the strategies of the model owner $S$, the data owners $D_{n}(n=1,\ldots,N)$, and the computing centers $C_{m}(m=1,\ldots,M)$. These strategies are taken in order: first, the model owner $S$ determines the optimal total model payment $\eta^{*}$; then, each data owner $D_{n}$ decides the quantity and quality of data to provide, i.e., the model quality contribution $q_{n}^{*}$; finally, each computing center $C_{m}$ determines the beneficial matching relationship $G^{*}_{m}$ with the specific data owner based on the optimal quantity of data undertaking $d^{*}_{m}$. To protect data privacy, we perform a one-time matching between data owners and the computing centers. In this manner, we formalize the resource-decoupled FL strategy as a Tripartite Stackelberg Model in Def. 1, where the model owner acts as the leader, the data owners act as secondary leaders, and the computing centers act as followers. The Stackelberg model establishes the sequential achievement of sub-utility optimization goals based on the major-minor relationships between the three parties. Within this framework, $N$ data owners and $M$ computing centers adopt strategies in their inner Nash game, ultimately achieving global utility optimization.
@@
The game comprises the strategies of the model owner $S$, the data owners $D_{n}(n=1,\ldots,N)$, and the computing centers $C_{m}(m=1,\ldots,M)$ to achieve global utility optimization in resource-decoupled FL.
(8) $\left\{\begin{array}[]{ll}\text{Leader }S:\quad\eta^{*}=\arg\max\limits_{\eta}U_{s},\\[5.0pt] \text{Sub-Leaders }D_{n}:\quad q_{n}^{*}=\arg\max\limits_{q_{n}}U_{n},&n=1,\dots,N,\\[5.0pt] \text{Followers }C_{m}:\quad G^{*}_{m}=\arg\max\limits_{G_{m}}U_{m},&m=1,\dots,M.\end{array}\right.$
We aim to generate the strategy profile $\left\langle\eta^{*},Q^{*},G^{*}\right\rangle$ of the tripartite Stackelberg model defined by Equ. 8 to achieve the SNE. Specifically, for the sub-game between data owners, we derive its corresponding Nash equilibrium (Appendix A.1), and the decisions of the three parties collectively achieve the SNE. We define SNE in the resource-decoupled FL, as shown in Equ. 9.
@@
(9) $\left\{\begin{array}[]{ll}U_{s}(\eta^{*},Q^{*},G^{*})\geq U_{s}(\eta,Q^{*},G^{*}),\\[8.0pt] U_{n}(\eta^{*},q_{n}^{*},q_{-n}^{*},G^{*})\geq U_{n}(\eta^{*},q_{n},q_{-n}^{*},G^{*}),&n=1,\ldots,N,\\[8.0pt] U_{m}(\eta^{*},Q^{*},G_{m}^{*})\geq U_{m}(\eta^{*},Q^{*},G_{m}),&m=1,\ldots,M.\end{array}\right.$
where $Q^{*}=\{q^{*}_{1},q^{*}_{2},...,q^{*}_{N}\}$ is the Nash equilibrium solutions of sub-games. SNE indicates that each of the three parties takes its optimal strategy, which maximizes its utility. No one can add its own utility by unilaterally changing its strategy.
@@ 3.3. Analysis of Tripartite Stackelberg Model
The policy space of the computing center is the quantity of data undertaken. This interval is $[0,|d_{m}|]$, which is bounded, closed, and continuous. Based on the utility function $U_{m}$, we obtain the first and second derivatives of $U_{m}$.
@@
For every computing center there exists a $d_{m}^{*}$ such that $U_{m}$ is optimal, we can also prove that there exists a unique optimal Nash equilibrium strategy, and the proof is similar to Lemma 4. Accordingly, the derived $d_{m}^{*}$ is given in Equ. 12.
(12) $d_{m}^{*}=\lambda\frac{(M-1)\sum_{n=1}^{N}\rho x_{n}}{\varepsilon\sum_{i=1}^{M}{\sigma_{i}}}(1-\sigma_{m}\frac{M-1}{\sum_{i=1}^{M}{\sigma_{i}}})$
@@
For data owner $q_{n}$, there is a unique Nash equilibrium optimal strategy $q_{n}^{*}$ to maximize its $U_{n}$ for the given $\eta$ and $q_{-n}$.
@@
The policy space for the data owner is to decide how much model quality to contribute. This interval is $[0,|Xn|]$, which is bounded, closed, and continuous. According to the utility function $U_{n}$ of the data owner, we obtain the first and second-order derivatives of $U_{n}$ as shown in Equ. 13 and Equ. 14.
@@
For the model owner, there is a unique optimal $\eta^{*}$ to optimize the utility $U_{s}$ for given the parameter of $\alpha$ and the objective data quality provided by the data owner $f_{n},n\in(0,N]$.
@@
The range of $\eta$ is $[0,\infty]$ for each model owner. Thus, this set of strategies is clearly convex and continuous. Although the value of $\eta$ can be infinite in theory, practically, the model owner will not give an infinite total model payment, and there is also the effect of diminishing marginal utility. This leads to the fact that $\eta$ is necessarily bounded and closed. We obtain the first-order and second-order derivatives as shown in Equ. 16 and Equ. 17.
@@
where $T_{n}=\frac{(N-1)}{\lambda\rho\sum_{i=1}^{N}{\frac{1}{f_{i}}}}(1-\frac{N-1}{f_{n}\sum_{i=1}^{N}{\frac{1}{f_{i}}}})$. Since $g(\cdot)$ is a convex function and to simulate the diminishing marginal benefit, we set $g(x)=ln(1+x)$ in this work. ∎
@@
See Appendix A.2. ∎
@@ 3.4. QD-RDFL
In this section, we design the QD-RDFL algorithm to identify the optimal strategy profile, in which a dynamic optimization mechanism is designed to improve the optimal strategy by evaluating the model contribution of data owners.
Using the backward induction method(Li et al., 2022; Kicsiny et al., 2014), we first calculate the optimal strategy of the model owner $\eta^{*}$ according to Equ. 18. We then calculate the optimal strategy $Q^{*}_{n}$ for $D_{n}$ according to Equ.14. Since $Q_{n}$ is evaluated along two dimensions, data quality is a posterior metric the server assesses after training, and data quantity $X^{*}_{n}$ is a decision determined immediately. Therefore, we initialize $D_{n}$ with an initial $f_{n}$ for computation and dynamically adjust $f_{n}$ based on the real training results. Finally, we calculate the optimal $d^{*}_{m}$ according to Equ. 12, and construct the optimal match $G^{*}$ using the Gale-Shapley algorithm (Teo et al., 2001; Gokhale et al., 2024). The matching remains unchanged during subsequent computations for private protection, which means adjusting $f_{n}$ will not affect the matching relationship.
To evaluate the real contribution of data quality, the most accurate method is to evaluate the test accuracy of each local model. However, it incurs additional costs and resource consumption. In contrast, accessing training loss does not require extra overhead, as it is generated during training. Therefore, we evaluate $f_{n}$ using the training loss during each iteration as shown in Equ. 19.
@@
where $\text{loss}(t)$ denotes the loss in a specific round $t$, $t_{s}$ represents the start time of validation, and $t_{e}$ represents the end time. Clients must submit their local model updates within the specified time range; otherwise, the edge server will reject the submission.
@@
0:  $D_{n}$ ($n=1,\ldots,N$), and $C_{m}$ ($m=1,\ldots,M$), $\rho$, $\varepsilon$, $\alpha$, $\lambda$, and adjusting round $L$
0:  the optimal strategy profile ¡$\eta^{*},Q^{*},G^{*}$¿ for model owner, data owners, and computing centers
1:  Initialization of $S$, $D$, $C$, report initial $f_{n}$
2:  Calculate $\eta^{*}$ // according to Equ. 18
3:  Calculate $x_{n}^{*}$ for $D_{n},n\in N$ // according to Equ. 15
4:  Calculate $d_{m}^{*}$ for $C_{m},m\in M$ // according to Equ. 12
5:  $G^{*}=Gale-Shapley(D_{n},C_{m})$ // call Gale-Shapley
6:  for each $D_{n},n\in[1,N]$ do
7:     Randomly pick $x_{n}^{*}$ and secure data transfer according to $G^{*}$
8:  end for
9:  for each round $t=1,2,...,T$ do
10:     for each $C_{m},m\in[1,M]$ do
11:        $w^{t+1}_{m},f^{t+1}_{m}\leftarrow\text{ClientUpdate}(C_{m},w^{t})$
@@
13:     if $t=L$ then
14:        Normalized each $f_{m}$ and recalculate $\eta^{*}$ and $x^{*}_{n}$
@@
16:           Pay additional training fees
@@
18:     end if
19:     $w^{t+1}\leftarrow\frac{1}{M}\sum^{M}_{m=1}w^{t+1}_{m}$ // global model aggregation
20:  end for
@@
0:  computing center $C_{m}$, global model parameter $w$, start time $t_{s}$, and end time $t_{e}$
0:  local model parameter $w_{m}$, data quality contribution $f_{m}$
@@
6:  $f_{m}=loss(t_{s})-loss(t_{e})$
7:  return  $w_{m}$, $f_{m}$
Consequently, the steps of QD-RDFL are outlined as follows, and the details are presented in Alg. 1 and Alg. 2.
@@
  • Step 2: Initial Optimal Strategy. $S$, $D$, and $C$ compute the optimal strategy profile $\langle\eta^{*},Q^{*},d^{*}\rangle$ for the first time based on Equ. 8 and Equ.9.
  • Step 3: Gale-Shapley Matching Data and Computing Resources. $D$ is matched with $C$ according to the initial optimal strategy, ensuring secure data transfer.
@@
In step 3 (line 5), we call the Gale-Shapley algorithm detailed in Appendix A.3, which performs an optimal matching $G^{*}$ based on the calculated $d^{*}_{m}$ as preference. In Step 5, after the data quality $f_{m}$ is adjusted according to the training results (line 11), the model owner and the data owners sequentially adjust $\eta^{*}$ and $x^{*}_{n}$ at the specific training round $L$ (line 14). $L$ is set to the number of rounds that best reflects the quality of the model, depending on the dataset. Subsequently, data owners providing more data pay additional training fees to their corresponding computing centers.
@@ 4. Experiments
Q4: How do different components (dynamic adjustment mechanisms) affect the global model performance?
@@ 4.1. Experiments Setting
We conduct extensive experiments on various vision datasets, including MNIST(LeCun and Cortes, 2005), CIFAR-10(Ayi and El-Sharkawy, 2020), and CIFAR-100(Krizhevsky, 2009). To simulate the variation in data quality provided by different data owners, we add Gaussian noise with varying intensity to the data from different owners. We calculate the mean square error (MSE) of the images after and before noise addition and use 1-MSE to represent the initial data quality of the data owner. The actual $f_{n}$ is evaluated using Equ. 19. Furthermore, to represent the relative position of $f_{n}$, we normalize its value. Other detailed settings of the experiment see Appendix A.4. The matching results of the data centers, the utility, and the parameter analysis are detailed in Appendix A.6. Our code is available at https://anonymous.4open.science/r/QD-RDFL-4B3F.
@@ 4.2. Experimental Results
SNE verification (RQ1). To confirm the existence and uniqueness of the SNE in QD-RDFL, we set $N=10$ on the MNIST dataset and observe the variation of utility under different strategies of the model owner and the data owner, respectively.
The uniqueness of the model owner’s utility $U_{s}$ is demonstrated in Fig. LABEL:sub@fig2a. We set the quantity and the strategies of the data owners remain constant, the curve of $U_{s}$ is convex with increasing $\eta$, with its highest point representing the optimal response. The optimal $\eta$ is 1.33 at the Stackelberg equilibrium, and the corresponding $U_{S}$ is about 0.57. The uniqueness of the Nash equilibrium among the data owners is demonstrated in Fig. LABEL:sub@fig2b. For the selected data owner $D_{n}$, we retain $x_{-n}$, the initial $f_{N},N=10$ and $\eta$, and then vary $x_{n}$. The results show that $U_{n}$ is convex, the maximum $U_{n}$ of 0.0066 is obtained when $x_{n}$ is 0.087. The results indicate that there are unique SNE in our method. Each data owner can maximize utility by determining the data contribution strategy in our method.
@@
Figure 3. Effect of initial $f_{n}$ to $U_{s}$ and the accuracy of global model on MNIST, (a) comparison of $U_{s}$ with different proportions of data owners and different report deviation ratio $f_{n}$, and (b) the accuracy of global model with different $f_{n}$.
Effect of initial $f_{n}$ (RQ2). To observe the effect of the initial $f_{n}$ on the utility of the data owner and the utility of the model owner, we design the following two sets of experiments.
For data owner: We select several data owners from $N,N=10$ for the experiment. For each chosen data owner, we fix the $f_{n}$ of the other data owners and increase or decrease the selected data owner’s $f_{n}$. We then observe the variation of $U_{n}$. For model owner: We increase or decrease all given data owner’s $f_{n}$ and observe the resulting variation of $U_{s}$.
For selected data owners, each utility $U_{n}$ increases as the deviation ratio of reported $f_{n}$ increases from -60% to +60% in the MNIST and CIFAR-10 datasets. For more details of the experiment, see Fig. 7 in Appendix A.5. As shown, the results suggest that the personal utility of the data owner $U_{n}$ is sensitive to data quality contribution $f_{n}$. $U_{n}$ increases monotonically with the reported $f_{n}$. Therefore, a valid evaluation of the data quality contributions in training processes is reasonable and necessary since it prevents data owners from falsely reporting quality contributions to get additional utility.
@@
Figure 6. The global model performance achieved by QD-RDFL in each round on different datasets
@@
Effect of QD-RDFL (RQ3). To verify the incentive effectiveness of our QD-RDFL, we perform comparison experiments with the following baselines on three different datasets.
@@
  • Random $\eta$: The model owner publishes an FL training task with a random payment value, and all users determine their strategies based on this payment.
@@
Varying numbers of data owners, we observe the utility of model owners $U_{s}$ and the average utility of $N$ data owners $U_{n},N\in N$, respectively. As shown in Figs. LABEL:sub@fig4a, LABEL:sub@fig4b and LABEL:sub@fig4c, $U_{s}$ of QD-RDFL always make outstanding performance compared to the benchmarks compared to the benchmarks on different datasets. The results indicate that the QD-RDFL algorithm can encourage data owners to provide more quantity and higher quality data for FL, thus getting the optimal utility of the model owner. However, due to the diminishing marginal effect, the increased rate of $U_{s}$ gradually slows down. Figs. LABEL:sub@fig5a, LABEL:sub@fig5b and LABEL:sub@fig5c show the average of $U_{n}$ on the different datasets. As $N$ increases, the average of $U_{n}$ decreases. Because as $N$ increases, more data owners participate in FL training while the total model payment grows relatively slowly. As shown, the average of $U_{n}$ of our QD-RDFL always outperforms compared to the benchmarks. The results demonstrate that data owners can always get better utility, thus effectively encouraging them. It is worth noting that since each set of experiments generates an $\eta$ independently for the Random algorithm, the results always oscillate, but it is always lower than our method.
@@
Ablation study (RQ4). To evaluate the effectiveness of the proposed dynamic adjustment mechanism, we verify accuracy of the global model on different datasets ($N=10$) and design the following ablation experiment, in which QD-RDFL without Dynamic Adjustment is set. In QD-RDFL, three parties first determine the initial strategies based on the data owners reported $f_{n}$. Then, QD-RDFL evaluates the actual $f_{n}$ in the specific training round ($L=3$) and dynamically adjusts the strategies. QD-RDFL without Dynamic Adjustment means that the reported $f_{n}$ is always used.
In Fig. 6, we verify the accuracy of QD-RDFL and QD-RDFL without Dynamic Adjustment over 70 training iterations on different datasets. The results show that the accuracy of QD-RDFL is better than that of QD-RDFL without Dynamic Adjustment. It illustrates that the dynamic adjustment mechanism performs a more accurate evaluation of data quality, thereby successfully encouraging data owners to provide high-quality and large-quantity data and improving the performance of the global model.
5. Conclusions and Future works